OpenStudy (natalie1234):

WILL GIVE MEDAL AND FAN FOR BEST ANSWER 1.arrange these numbers from least to greatest 5/6, 7/9, 3/5 2.Arrange the list of numbers in increasing order. 2/5, 4/9, 3/7 3. Which of these numbers is between 1 and 3/4 4.Write 7/20 as a percent and a decimal.

OpenStudy (anonymous):

The easiest to order them is to convert to decimal form.

OpenStudy (anonymous):

divide the numerator by the denominator = the decimal form

OpenStudy (natalie1234):

oh ok

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

5/6= .83 7/9=.87 3/5= .6

OpenStudy (anonymous):

then compare then write them in the form.

OpenStudy (anonymous):

so it would be 3/5 , 5/6 , 7/9 right?

OpenStudy (natalie1234):

yeah (:

OpenStudy (anonymous):

increasing order i am assuming would be least to greatest right?

OpenStudy (anonymous):

the seconed one should probly be 3/5 don't you think

OpenStudy (natalie1234):

yeah

OpenStudy (anonymous):

u wanna the the second urself?

OpenStudy (anonymous):

no it shouldnt

OpenStudy (anonymous):

because .60 is smaller than .83

OpenStudy (natalie1234):

my options for number 2 are 2/5, 3/7, 4/9 -3/7, -7/9, -3/5, -5/6 -7/9,-3/7,-5/6,-3/5

OpenStudy (natalie1234):

oops i missed one the other is -3/7, -3/5,-7/9,-5/6

OpenStudy (anonymous):

2/5 , 3/7 , 2/9 is it right?

OpenStudy (anonymous):

i mean 4/9

OpenStudy (natalie1234):

yeah thanks for your help i got the other questions (:

OpenStudy (anonymous):

kk ur welcome.. what grade is this?